What is the value of a share of stock of HOV Inc. to an investor who requires a 12 percent rate of return if HOV's current dividend is $1.20? Assume earnings and dividends are expected to grow at a compound annual rate of 7 percent.
Answer #1
Value of a share using the constant dividend growth model = D0*(1+g)/(r-g)
where
D0 = Current dividend
g = growth rate in dividends
r = required return
Substituting available values we have
Value of the share = 1.20*1.07/(0.12-0.07) = $    25.68